OpenStudy (anonymous):

Business Education Nursing Computer Science Other Total Men 38 26 14 47 18 143 Women 23 39 26 32 28 148 Total 61 65 40 79 46 291 If a student is randomly chosen, what is the probability that the student is a business or nursing major? Question 9 options: 0.0289 0.0288 0.347 0.694

OpenStudy (kira_yamato):

P(Business) = 61/291 P(Nursing) = 79/291 P(Business or nursing) = (61+79)/291 = 140/291